OpenStudy (anonymous):

I need to find S25 for -7+-2+3+8+...

OpenStudy (anonymous):

@eric_d

OpenStudy (solomonzelman):

So you have a difference of 5. \(\large\color{darkblue}{ \rm (-7) +5(25-1)=a_{25}}\)

OpenStudy (solomonzelman):

Can you find the \(\large\color{darkblue}{ \rm a_{25} ~}\) ?

OpenStudy (anonymous):

-7+5(24)=a25 -2(24)=a25?

OpenStudy (anonymous):

-48=a25?

OpenStudy (solomonzelman):

your second step is incorrect.

OpenStudy (solomonzelman):

\(\large\color{darkblue}{ \rm (-7) +5(25-1)=a_{25}}\) \(\large\color{darkblue}{ \rm (-7) +5(24)=a_{25}}\) \(\large\color{darkblue}{ \rm (-7) +120 =a_{25}}\) \(\large\color{darkblue}{ \rm 113 =a_{25}}\)